If a household’s income falls from R12 000 to R10 000, and its consumption falls
from R9 500 to R8 000, then:
(2)
(1) The marginal propensity to consume is ‐0.8.
(2) The marginal propensity to consume is 0.75.
(3) The marginal propensity to consume is 0.2.
(4) The marginal propensity to save is 0.15.
1
Expert's answer
2020-09-23T10:21:15-0400

marginal propensity consume "=" "\frac{change in consumption}{change in income}"


"= \frac{8000-9500}{1000-12000}"

"=\frac{-1500}{-2000}"

"=" "0.75"

so option 2 is correct